OpenStudy (anonymous):

Solve the intitial value problem: dy/dx=x(6+x^2)^6, y(0)=0

OpenStudy (turingtest):

just separate the variables and integrate

OpenStudy (turingtest):

\[\frac{dy}{dx}=x(6+x^2)^6\]\[dy=x(6+x^2)^6dx\]\[\int dy=\int x(6+x^2)^6dx\]what do you get? (don't forget the integration constant for the indefinite integral)

OpenStudy (turingtest):

...the integral involves a u-substitution

OpenStudy (anonymous):

[(x^2+6)^7]/14

OpenStudy (anonymous):

So...did I do that right?

OpenStudy (dumbcow):

yes good job now use initial condition y(0) = 0 to solve for the constant (x^2+6)^7 /14 + C = 0 .....x=0

OpenStudy (anonymous):

Thank you. I got y=1/14(6+x^2)^7-(139968/7) I feel confident with that

OpenStudy (dumbcow):

as you should :)